OpenStudy (anonymous):

Solve the following system of equations algebraically. Verify your solution either graphically or by using matrices 3x-y=0 5x+2y=22

OpenStudy (nurali):

3x-y=0 5x+2y=22 y = 3x 5x + 2(3x) = 22 5x + 6x = 22 11x = 22 x = 2 y = 3(2) y = 6 (x,y) = (2,6)

OpenStudy (nurali):

Verify Equation 3x-y=0 3(2)-6=0 ... Plug in (2,6) 6 - 6 = 0 0 = 0 So (2,6) works for the first equation ------------------------------------ 5x+2y=22 5(2)+2(6)=22 ... Plug in (2,6) 10+12 = 22 22 = 22 and (2,6) works for the second equation

OpenStudy (anonymous):

okay so 2,6 is the answer is for both equations, i dont understand how to verify it though

OpenStudy (anonymous):

so i plug it in but i'm not getting it

OpenStudy (anonymous):

3x-y=0 3(2)-6=0 ... Plug in (2,6) 6 - 6 = 0 0 = 0 that's how you verify it? so i plug in 2,6 and it has to come out to 0?

OpenStudy (nurali):

if the answer 0 means it is verify the value (2,6) is correct.

OpenStudy (nurali):

Support if x=3 and y=6 put in value 3(3)-6=0 9-6=0 3=0 it means the x value is not verify so the value of x is wrong.

OpenStudy (anonymous):

ok so i should put 3x-y=0 3(2)-6=0 6 - 6 = 0 0 = 0 to show thats how you verify it? and put (x,y) = (2,6) as the answer to the equation?

OpenStudy (nurali):

Yes